# KinesCeTI: A Modular and Size-Adaptable Force Feedback Glove with Interchangeable Actuation for the Index and Thumb

### v1

@@ Abstract
Force feedback gloves in haptic applications remain constrained by limited adaptability, simplified feedback and fixed architectures that limit force feedback versatility. To address these challenges, we present KinesCeTI, a modular force feedback exoskeleton for the index and thumb, designed as a multipurpose device adaptable to a wide range of hand sizes. The glove incorporates interchangeable thimbles for fingertip or phalanx attachment and a bidirectional tendon transmission that supports both passive and active feedback. It is combined with a modular actuation design, where different feedback systems may be attached. The system was tested with two actuation modules: a compliant ratchet-pawl braking mechanism for passive feedback and a novel one-way clutch for variable active feedback, newly introduced here. The system was evaluated in three user studies with 20 participants each, assessing ergonomics, actuation performance and usability in both real and virtual tasks. Results indicate that the glove adapts to different hand sizes and provides effective feedback with both mechanisms, highlighting its potential as a versatile platform for haptic research.
footnotetext: This manuscript was submitted to IEEE Transactions on Haptics on October 8, 2025. It has not been peer reviewed. This preprint may differ from the final published version.
@@ I Introduction
Haptic technology has received growing attention in recent decades, following the development of acoustic and visual interfaces throughout the previous century. As their performance reached advanced levels, the integration of touch as an additional modality emerged as a frontier in human-machine interaction. However, despite its progress in recent years [yang2021recent], realistic haptic feedback remains challenging due to technological limits, the complexity of touch and application demands. Haptic systems have been explored in multiple areas, including medicine [licona2019applications], cultural heritage [brewster2017impact], affective touch [eid2015affective], teleoperation [pacchierotti2023cutaneous] and Augmented and Virtual Reality (AR/VR) [tong2023survey]. Its relevance is highlighted in concepts such as the Tactile Internet [fettweis20145g] and the Metaverse [lee2024all].
Because skin sensitivity varies across the body [myles2007tactile], haptic designs are heterogeneous [pacchierotti2017wearable, basdogan2020review, yin2021wearable, culbertson2018haptics, giri2021application]. However, out of all body parts, the hands are the main target for haptic research, due to their central role in manipulation and interaction tasks. Due to their dexterity, sensitivity and force capabilities, numerous types of devices have been developed. These include thimbles, gloves and externally grounded interfaces, each exploring diverse type of stimuli, including pressure, force and vibration [pacchierotti2017wearable]. Nevertheless, no standard design has emerged yet.
This work focuses on a specific category of haptic interfaces: kinesthetic or force feedback gloves. These systems have been extensively researched and developed, with various designs exploring different actuation and sensing technologies, transmission systems and overall structures [tiboni2022soft, du2021review, ferguson2020hand, caeiro2021systematic]. However, as discussed in the following section, current solutions face limitations in feedback capabilities, adaptability to different users and ergonomics.
This paper outlines key design considerations and presents a modular and adaptive hand exoskeleton that implements them. Its actuation potential is demonstrated with two customized force feedback mechanisms developed in this work, which are evaluated alongside the system’s ergonomics and mobility in three user studies. The results are discussed in relation to the glove’s features compared to the state of the art and its potential as a platform for haptics research.
@@ II Force Feedback Gloves: Design Considerations and State of the Art
The design of haptic gloves must begin with the human hand: its size, strength and mobility. Anthropometric studies for healthy subjects report large variability in hand and finger dimensions [chen2011human, mclain2010use, tilley2001measure]. The hand is also capable of considerable force, with mean grip forces up to 55-56 kg and pinch forces around 10 kg [mathiowetz1985grip, angst2010prediction]. Regarding mobility, the human hand has 21 Degrees of Freedom (DOFs) in the fingers and thumb [wang2018toward], which are distributed as follows. Each finger (index, middle, ring, little) includes three joints (metacarpophalangeal (MCP), proximal interphalangeal (PIP) and distal interphalangeal (DIP)) with flexion/extension at every joint and abduction/adduction at the MCP, totaling four DOFs. The thumb consists of three joints (carpometacarpal (CMC), MCP and interphalangeal (IP)) with flexion/extension at each of them and abduction/adduction at the CMC and MCP, totaling five DOFs. Taking into account the hand’s capabilities, the ideal haptic glove would therefore include: high-fidelity force feedback with low latency, accurate motion tracking below perceptual thresholds, ergonomic design, imperceptibility when inactive, adaptability across hand sizes, ease of use, safety and affordability.
From a design perspective, each desired feature imposes new requirements that are often interrelated and sometimes contradictory. To analyze them, we adopt the architecture described by Wang et al. [wang2018toward], which divides exoskeletons into five subsystems: sensing, actuation, transmission, control and structure. Design possibilities for each subsystem are extensive [du2021review, sarac2019design], although interdependent. For instance, the type of control is constrained by sensing and actuation.
Within this framework, we can perform an evaluation of the ideal characteristics, being the most restrictive aspects motion detection and haptics. With regards to feedback modalities, from this point forward this work focuses exclusively on kinesthetic feedback, leaving tactile and thermal modalities outside the scope of this work. Even with this limited approach, motion tracking and force feedback must address the multiple DOFs of the hand. In the context of glove design, DOFs can be categorized in three levels. The first type are passive DOFs, the simplest to implement, as they only imply passive motion. Sensed DOFs add complexity by also measuring movement, requiring electronics and, when needed, markers or attachment elements. And last, feedback DOFs, the most challenging in terms of design, requiring the implementation of actuators, their corresponding force transmission and attachment systems and additional control and electronics. Each type entails increasing demands in terms of weight, volume, power and complexity.
Ideally, every DOF would be both sensed and susceptible to feedback. In practice, the limited space on the hand, the high number of DOFs and technological constraints (particularly in actuation) require trade-offs. Moreover, actuated DOFs are not inherently a binary feature (force vs no-force) but involve various qualitative dimensions, required for achieving complete kinesthetic feedback. Some authors, like Wang et al. [wang2018toward], have analyzed such aspects. In this work, we present an expanded version of different force feedback dimensions:
@@
  • Feedback direction, being some systems capable of feedback in only one direction for a single DOF (e.g., resistance during hand closing), while the ideal system would enable bidirectional feedback. Moreover, bidirectional systems may provide force in both directions simultaneously or at either of them at a particular moment.
  • Transition between states. Being the actuation system capable of switching seamlessly and immediately between the free motion and the feedback states. Such feature requires backdrivability, in order to experience minimal resistance when no force is provided, while enabling immediate feedback when required.
  • Stiffness generation. To recreate different types of objects and interactions, the system should render a wide range of stiffness profiles, from soft objects (e.g., foam) to rigid boundaries (e.g., a glass surface).
  • Type of force. Systems may generate passive force, which resist motion (e.g., friction and brake mechanisms), or active forces, that in addition can induce movement. The latter offers a broader range of feedback capabilities, although at the expense of increasing complexity.
@@
Meeting both these dimensions and the aforementioned overall requirements is highly challenging, since all glove subsystems must support them and many are mutually conflicting. High-fidelity tracking and feedback across all 21 DOFs demand extensive actuation and sensing, yet current technologies make such designs heavy, bulky and power-intensive, reducing wearability and ergonomics. As a general rule, simpler unidirectional passive systems are easier to implement, while bidirectional active designs are far more complex. A further trade-off exists between anatomical adaptability and structural design: soft gloves provide good ergonomics but limited user adjustability, whereas rigid exoskeletons can adapt to different hand sizes at the cost of comfort.
@@
While numerous design strategies have been proposed [pacchierotti2017wearable, du2021review, wang2018toward], all force feedback gloves inevitably involve trade-offs, which are best illustrated by the most representative commercial systems. The leading solutions are HaptX G1 [haptx_glove] and Senseglove Nova 2 [senseglove] with soft-glove architectures. Both provide tactile and force feedback, using a tendon based force transmission system to restrict the grasping motion. This approach yields compact and lightweight designs but limits force feedback to one direction. Their actuation technologies differ, with SenseGlove employing magnetic friction brakes, while HaptX uses a pneumatic brake system, delivering in both cases only passive resistance, complemented by tactile cues for immersion. For hand-size adaptability, both rely on different glove sizes.
An alternative to soft gloves are linkage exoskeletons, which can potentially provide bidirectional feedback and adapt to different users at the cost of higher weight and volume. Dexmo is a prominent example, being originally passive [gu2016dexmo] while in its latest version generates active feedback using customized servomotors. SenseGlove’s earlier DK1 also followed a linkage-based structure, but provided exclusively unidirectional feedback. Numerous other prototypes exist, such as Grabity [choi2017grabity], which simulated weights and opposition grip forces and Wolverine [choi2016wolverine], capable of providing high braking forces with low power consumption. These, however, have not been further developed.
@@
Across these examples, several common limitations of current force feedback gloves emerge. Simplified and often unidirectional feedback is favored to preserve compactness and ergonomics. Hand adaptability is typically addressed through detachable gloves or multiple sizes, but in many devices remains limited. Furthermore, the majority of these solutions adopt closed architectures tailored to a single application or feedback profile, reducing versatility and generalizability. An additional overlooked factor is the acoustic perception of the system. Force feedback presents inherent actuator noise that may influence how users perceive both the feedback and the overall interaction and should therefore be considered in the design process.
@@
These limitations highlight the need for a versatile, modular glove architecture that supports multiple feedback modalities, adapts to different hand sizes and explicitly considers acoustic performance. The following section presents such an exoskeleton, designed with interchangeable actuation and thimble modules and a structural design adaptable to different users.
III KinesCeTI: System Design and Actuation Mechanisms
@@ III-A Design goals
The previous section reviewed design aspects and highlighted the limitations of current technologies and approaches. Building on this analysis, we present the KinesCeTI exoskeleton, developed to address these challenges. It’s designed based on the following main goals:
  • Adaptable to a wide range of adult hand sizes.
  • Unhindered hand motion with high backdrivability, enabling force feedback only when required.
  • General purpose force feedback device for haptic applications, with especial focus on AR/VR. Thus, must be possible to implement different types of force feedback.
@@ III-B Overall structure
After establishing these design decisions, we proceed to choose the subsystems that enable them. The first consideration is the type of structure and transmission system. Although soft actuation systems have advanced significantly in the last decade, they still lack the maturity and accessibility required for bidirectional systems [tiboni2022soft]. Consequently, a linkage-based architecture placed on the dorsal of the hand is chosen, being both wearable and portable. Because it is grounded to the hand itself, kinesthetic feedback can be applied only to the fingers relative to the hand, but not to the hand as a whole.
One important design decision is the manufacturing method. To ensure low cost and accessibility, most components are made of common polymer-based materials (PLA, PETG, TPU), with the exception of commercial structural parts (e.g., bearings, bolts) and electronics. The exoskeleton is produced through rapid prototyping with additive manufacturing, specifically using Prusa MK3S+ and Prusa Mini+ 3D printers.
@@
In order to focus on developing and testing an effective design, this exoskeleton is designed for the index finger and the thumb, as together they represent all fundamental motion types of the human fingers. Since the index is anatomically similar to the middle, ring and little fingers, its design can be adapted to those in future iterations. In contrast the thumb has a unique anatomy and kinematics, requiring a different mechanical design.
@@
A linkage architecture is implemented for both fingers. The base structure consists of a rigid dorsal platform, attached to the hand with two adjustable strap bands with buckles: one around the knuckles and the other around the wrist. For ergonomic purposes, its contact area with the hand consists of a dorsal-shaped TPU structure with a layer of neoprene rubber for comfortable skin contact. The electronics hub, referred to as the Main Hub, is mounted on this platform. All actuators and sensors are connected to it throughout different connection modules distributed across the exoskeleton. For hand orientation tracking, a 9-DOF IMU (Adafruit BNO055) is embedded within the Main Hub.
@@
The index structure is adjacent to the Main Hub and incorporates a revolute joint that enables abduction/adduction of the index, with its angle measured by a contactless magnetic rotary encoder (AS5600), as shown in Fig. 6. Mounted above are the Knuckle Hub with the linkage and transmission systems, which are nearly identical for both index and thumb, differing only in linkage lengths (explained later). An overview of the exoskeleton is shown in Fig. 1. As the thumb’s ab/adduction produces a wide motion range, the base structure must adapt accordingly. To this end, a structural extension is designed for the thumb oriented at 120 from the index finger in the coronal plane and tilted 30 downward in the sagittal plane (plane conventions follow [bullock2012assessing]).
Figure 1: Overview of the KinesCeTI exoskeleton, showing dorsal placement, Main Hub and attachment system.
@@
To enable unrestricted motion of the thumb, the base structure incorporates two sensed and actuated DOFs. This system, shown in Fig. 2(c), consists of two concentric shafts, with their angles measured by two AS5600 on the outer structure. The first sensed DOF of rotation is for the overall shaft and referred to as simple ab/adduction, allowing up to 90 of motion. The second sensed DOF, combined ab/adduction, includes both this first motion and the ab/adduction of the Knuckle Hub on top of it, transmitted through a bevel gear mechanism. With this bevel mechanism, the thumb’s Knuckle Hub has a range of motion of $\pm 90^{\circ}$ (180 total), with a gear ratio of 13:19 between the Knuckle Hub and the main shaft.
All rotational systems have mechanical stops at the end of their motion range to ensure safe operation. Both shafts also receive feedback from customized ratchet-pawl mechanisms (Fig. 2(a)-(b)), described later in the actuation section. In addition, a compression spring applies a small pulling force to the mobile thumb structure when fully extended, reducing inertia and facilitating user motion.
@@
Figure 2: Thumb structure: (a) frontal view and ratchet-pawl mechanism for the bevel system, (b) rear view and ratchet-pawl mechanism for the simple ab/adduction shaft and (c) side view highlighting the main components.

### v2

@@ Abstract
Force feedback gloves in haptic applications remain constrained by limited adaptability, simplified feedback, and fixed architectures that limit force feedback versatility. To address these challenges, we present KinesCeTI, a modular force feedback exoskeleton for the index and thumb, designed as a multipurpose device adaptable to a wide range of hand sizes. The glove incorporates interchangeable thimbles for fingertip or phalanx attachment and a bidirectional tendon transmission that supports both passive and active feedback. It is combined with a modular actuation design, where different feedback systems may be attached. The system was tested with two actuation modules: a compliant ratchet-pawl braking mechanism for passive feedback and a novel one-way clutch for variable active feedback, newly introduced here. The system was evaluated in three user studies with 20 participants each, assessing ergonomics, actuation performance and usability in both real and virtual tasks. Results indicate that the glove adapts to different hand sizes and provides effective feedback with both mechanisms, highlighting its potential as a versatile platform for haptic research.
footnotetext: This manuscript has been accepted for publication in IEEE Transactions on Haptics following peer review. This version corresponds to the accepted manuscript and may differ from the final published version.
@@ I Introduction
Haptic technology has received growing attention in recent decades, becoming a key modality in human–machine interaction. However, despite progress [37], realistic haptic feedback remains challenging due to technological limits and the complexity of touch. Applied across diverse domains, haptic interfaces display highly heterogeneous designs [30, 14], with the hands as a primary focus of haptic research due to their central role in manipulation and interaction. Their capabilities have led to the development of a wide range of devices, including thimbles, gloves and externally grounded interfaces, which explore diverse stimulus modalities such as pressure, force, and vibration [30]. Nevertheless, no standard design has yet emerged.
This work focuses on kinesthetic or force feedback gloves, a class of haptic interfaces widely studied [33, 15, 8], although current solutions still face limitations in feedback capabilities, user adaptability and ergonomics. To address these challenges, key design considerations are outlined and a modular, adaptive hand exoskeleton is presented. Its actuation potential is demonstrated through two customized force feedback mechanisms developed in this work and evaluated together with the system’s ergonomics and mobility in three user studies. The results are discussed in the context of the state of the art and the system’s potential as a versatile haptics research platform.
@@ II Force Feedback Gloves: Design Considerations and State of the Art
The design of haptic gloves depends on the features of the human hand. Anthropometric studies report large variability in hand and finger dimensions [9, 17, 34], while the hand is capable of considerable force: mean grip forces up to 55 kg and pinch forces around 10 kg [4]. Regarding mobility, the digits exhibit 21 degrees of freedom (DOFs), with four DOFs per finger and five in the thumb [36]. Taking into account such capabilities, the ideal haptic glove would therefore include high-fidelity haptic feedback with low latency, accurate motion tracking below perceptual thresholds, ergonomic design, unobtrusive when inactive, adaptability across hand sizes, ease of use, safety, and affordability. From a design perspective, each desired feature imposes new requirements that are often interrelated and sometimes contradictory. To analyze them, we adopt the architecture described by Wang et al. [36], which divides exoskeletons into five subsystems: sensing, actuation, transmission, control, and structure. Design possibilities for each subsystem are extensive [15, 31] but interdependent, e.g. the type of control is constrained by sensing and actuation.
Within this framework, the most restrictive design aspects are motion tracking and haptics. Regarding feedback modalities, this work focuses exclusively on kinesthetic feedback and does not address tactile or thermal modalities. Even with this limited approach, motion tracking and force feedback must address the multiple DOFs of the hand. In the context of glove design, DOFs can be classified as passive, sensed and feedback DOFs. Each type entails increasing demands in terms of weight, volume, power and complexity. While ideally all DOFs would be sensed and actuated, practical constraints such as limited space, high DOF count, and technological constraints require trade-offs. Moreover, actuated DOFs are not inherently a binary feature (force vs no-force) but involve various qualitative dimensions for complete kinesthetic feedback. Building on prior analyses (Wang et al. [36]), this work presents an expanded set of force feedback dimensions:
@@
  • Feedback direction: some devices apply unidirectional force for a given DOF (e.g., resistance during hand closing), whereas an ideal system enables force in both directions, either alternately or simultaneously.
  • Transition between states: seamless and rapid switching between free motion and force feedback, with high backdrivability when no force is required.
  • Stiffness generation: rendering stiffness profiles ranging from compliant objects (e.g., foam) to rigid boundaries (e.g., glass surfaces).
  • Type of force: passive systems resisting motion (e.g., friction and braking mechanisms) versus active devices capable of inducing movement.
@@
Meeting these dimensions remains challenging, as many design requirements are mutually conflicting. High-fidelity tracking and feedback across all DOFs demand extensive actuation and sensing, yet current technologies lead to heavy, bulky and power-intensive designs that reduce wearability and ergonomics. In general, unidirectional passive systems are simpler, whereas bidirectional active designs are more complex. A further trade-off exists between anatomical adaptability and structural design: soft gloves offer better ergonomics but limited adjustability, whereas rigid exoskeletons can adapt to different hand sizes at the cost of weight, volume and comfort.
@@
Although numerous design strategies have been proposed [30, 15, 36, 23], force feedback gloves inevitably involve trade-offs, exemplified by commercial systems such as HaptX G1 [20] and Senseglove Nova 2 [32]. Both adopt soft-glove architectures, providing force feedback through tendon-based transmissions that restrict grasping motion, enabling compact and lightweight designs but limiting force feedback to one direction. SenseGlove and HaptX rely on braking-based actuation, using magnetic friction and pneumatic braking, respectively, to provide passive resistance. Hand-size adaptability is achieved through different glove sizes.
Linkage-based exoskeletons represent an alternative design approach, offering the potential for bidirectional force feedback and cross-user adaptability through an adjustable structure. Dexmo is a prominent example, originally passive [19] but later extended to active feedback using customized servomotors. SenseGlove’s earlier DK1 also adopted a linkage-based structure, but provided exclusively unidirectional feedback. Beyond fixed-design gloves, Bartalucci et al. [6] present a bidirectional, hand-size-adaptable glove using a modular architecture with magnetic couplings. However, actuation is offloaded to an external station via Bowden cables, limiting mobility, while the wearable structure weighs approximately 400 g. Chinello et al. [10] also explore modularity through separate kinesthetic and tactile modules. Finger-size adaptability is achieved through a sliding linkage, although kinesthetic feedback is limited to interphalangeal forces. Linkage-based exoskeletons are common in rehabilitation contexts, as exemplified by Iqbal et al. [24, 22], who implement active bidirectional feedback in direct-drive configurations, and linkage designs adaptable to different hands. In these designs, low backdrivability during free motion is compensated through active control, at the expense of increased power consumption. Conversely, Leonardis et al.[27] present a full upper-limb exoskeleton with haptic thimbles for teleoperation, offering extensive haptic capabilities at the cost of a large, complex, and stationary system with limited portability and wearability. At the opposite end of the design spectrum, some rigid designs prioritize palm-grounded, task-specific feedback, such as Grabity [11], which simulates weight and opposition grip forces, and Wolverine [12], capable of delivering high braking forces with low power consumption.
@@
Across these examples, common limitations of current force feedback gloves emerge. Existing systems tend to fall into two categories: simplified, often unidirectional designs that preserve compactness and ergonomics, or complex and heavy systems that lack adaptability or versatility. Hand-size adaptability is typically addressed through detachable gloves, multiple sizes or, in linkage-based designs, structures with limited adaptability that require prior adjustment or support attachment only at a single point. Moreover, most solutions adopt closed architectures tailored to a single application or feedback profile, reducing versatility and generalizability. Such closed architectures hinder research progress, as haptic researchers often need to develop an entire glove platform before exploring novel concepts. An additional and frequently overlooked factor is acoustic perception: force feedback inherently introduces actuator noise that may influence user perception of both the feedback and the overall interaction and should therefore be considered in the design process.
@@
These limitations highlight the need for a versatile, modular glove architecture that supports multiple force feedback dimensions, adapts to different hand sizes and considers acoustic perception. The following section presents such an exoskeleton, designed with interchangeable actuation and thimble modules and a structure passively adaptable to different users.
III KinesCeTI: Design and Actuation Mechanisms
@@ III-A Design goals
The previous section reviewed design aspects and highlighted the limitations of current technologies and approaches. Building on this analysis, we present the KinesCeTI exoskeleton (Fig. 1), designed around the following goals:
  • Adaptability across adult hand sizes, supporting attachment at either the fingertip or the second phalanx.
  • Passively unhindered hand motion with high backdrivability during free motion, ensuring minimal resistance in the absence of force feedback.
  • General purpose force feedback device for haptic applications, designed to enable different force feedback dimensions (uni-/bidirectional, active/passive).
  • Fully wearable and portable, excluding the need for external actuation stations.
  • Low-noise actuation to minimize acoustic annoyance.
@@ III-B Overall structure
Based on these design goals, the choice of subsystems that enable them is addressed, starting with the structural architecture. Although soft actuation systems have advanced significantly in the last decade, they still lack the maturity and accessibility required for bidirectional systems [33], as well as hand-size adaptability. A linkage-based architecture mounted on the dorsal side of the hand is therefore selected, enabling hand adaptability while being wearable and portable. As it’s grounded to the hand, kinesthetic feedback is limited to finger–hand interactions. Most components are fabricated from common polymer-based materials (PLA, PETG, TPU) to ensure lower costs and accessibility, with the exception of commercial elements such as bearings and electronic components. The exoskeleton is produced via additive manufacturing using Prusa MK3S+ and Prusa Mini+ 3D printers.
@@
To focus on developing an effective design, the exoskeleton targets the index finger and thumb, which together capture the fundamental motion types of the human hand. Since the index is anatomically similar to the middle, ring and little fingers, its design (after further optimizations) can be adapted to those in future iterations. In contrast, the thumb has a unique anatomy and kinematics, requiring a dedicated mechanical design.
@@
The selected architecture is implemented on a rigid dorsal platform attached to the hand via two adjustable strap bands with buckles positioned around the knuckles and the wrist. For ergonomic purposes, its contact area with the hand consists of a dorsal-shaped TPU structure combined with a neoprene rubber layer for comfortable skin contact. The electronics hub, referred to as the Main Hub, is mounted on this platform. All actuators and sensors are connected to it through different connection boards distributed across the exoskeleton.
@@
The index structure is adjacent to the Main Hub and incorporates a revolute joint that enables ab-/adduction, with its angle measured by a contactless magnetic rotary encoder (AS5600), as shown in Fig. 2. Mounted above is the Knuckle Hub with the linkage and transmission systems, which are nearly identical for both index and thumb, differing only in linkage lengths, as described in the following subsection. Due to the large motion range of ab-/adduction of the thumb, a structural extension is introduced, oriented at 120 from the index finger in the coronal plane and tilted 30 downward in the sagittal plane (plane conventions follow [7]).
Figure 2: Side view of the index assembly showing the linkage mechanism, force transmission system, and enabled DOFs.
@@
To enable unrestricted thumb motion, the base structure incorporates two sensed and actuated DOFs (Fig. 3(c)). The system consists of two concentric shafts, with angular positions measured by two AS5600 encoders mounted on the outer structure. The first DOF corresponds to rotation of the main shaft (simple ab-/adduction), allowing up to $90^{\circ}$ of motion. The second DOF (combined ab-/adduction) represents the superposition of this main shaft rotation and the ab-/adduction of the Knuckle Hub mounted above it, transmitted through a bevel gear mechanism. The Knuckle Hub itself allows an ab-/adduction range of $\pm 90^{\circ}$ (180 total), with a gear ratio of 13:19 between the Knuckle Hub and the main shaft.
@@
Figure 3: Thumb structure: (a) frontal view of the bevel system with ratchet–pawl mechanism, (b) rear view (housing transparent) of the simple ab-/adduction shaft with ratchet–pawl mechanism, and (c) cross-sectional view showing the main components.
All rotational systems have mechanical stops at the end of their motion range to ensure safe operation. Both shafts also receive feedback from customized ratchet-pawl mechanisms (Fig. 3(a)-(b)), described in the actuation subsection. A compression spring applies a small pulling force to the mobile thumb structure when fully extended, reducing inertia and facilitating user motion.